Circular No. 7486

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2000 S. N. Shore, Indiana University, South Bend; S. Starrfield, Arizona

State University; H. E. Bond and R. Downes, Space Telescope Science Institute; P. H. Hauschildt, University of Georgia; R. D. Gehrz and C. E. Woodward, University of Minnesota; J. Krautter, Landessternwarte, Heidelberg; and A. N. Evans, Keele University, write: “We observed this nova (cf. IAUC 7457, 7458) with the Hubble Space Telescope’s Imaging Spectrograph (STIS) on Aug. 19.7 and Aug. 20.9 UT using the medium-resolution échelle gratings (E140M and E230M). Three spectra were obtained, centered at 142.5, 197.8, and 270.7 nm with resolutions of ∼ 45 000 (E140M) and 30 000 (E230M). Weak iron-peak absorption may still be present, especially from 155.0 to 156.5 nm. The spectrum strongly resembles that of V382 Vel obtained 2 months after visual maximum (IAUC 7261) but with much stronger Lyα and Si III] 189.5-nm and C III] 191.0-nm emission, with Si III]/C III] ∼ 2. Strong C IV P-Cyg absorption is seen with a terminal velocity of 2000 km/s, similar to the FWHM velocity for the emission lines. The following strong emission lines have been detected: C III 107.6-nm, N V 124.0-nm, N III/O III 126.7-nm, O I 130.4-nm, C II 133.4-nm, Si IV/O IV] 140.0-nm, C IV 155.0-nm, He II 164.0-nm, O III] 166.7-nm, N III] 175.0nm, Si II 181.6-nm, Al III 186.0-nm, Si III] 189.5-nm, C III] 191.0-nm, N II] 214.5-nm, C II 232.1-nm, Al II 267.2-nm, and Mg II 280.0-nm. Unusual features of the spectrum are the absence of N IV] at 148.6 and 171.8 nm and the presence of strong Lyα emission at a time when He II 164.0-nm is also strong, but the Lyα in this nova is stronger due to lower interstellar absorption. Most emission profiles show fine structure consistent with knots of the sort detected in HST spectra of V1974 Cyg and V382 Vel. A strong interstellar absorption line spectrum is present, showing Si II/S II 126.0nm, Si IV 140.0-nm, C IV 155.0-nm, Al II 167.1-nm, Al III 186.0-nm, and Fe II 260.0-nm with Galactic and LMC components (+250 km/s) clearly resolved. The integrated flux from 115.0 to 312.0 nm was 8.6 × 10^{−11} erg cm^{−2} s^{−1} [or 5.6 × 10^{−10} erg cm^{−2} s^{−1}, corrected for LMC reddening with E(B − V) = 0.2]. For a distance of 50 kpc, this corresponds to 4.1 × 10^{4} L_{⊙}. In view of the similarity of the spectra, these data yield a probable distance for V382 Vel of 3000 pc, assuming its reddening is E(B − V) = 0.2. A further STIS observation is planned for mid-Nov., and continued groundbased monitoring at all wavelengths is urged.”

2000 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green